# U2UData: A Large-scale Cooperative Perception Dataset for Swarm UAVs Autonomous Flight

### v1

@@ 5.2. Cooperative 3D Object Tracking
When2Com(16, ) 24.30 43.68 60.69 52.62 36.75 26.21
DiscoNet(15, ) 24.58 44.71 62.79 55.70 39.81 23.89
V2VNet(12, ) 26.19 47.43 65.81 56.58 41.63 23.83
V2X-ViT(14, ) 26.85 46.50 64.50 55.92 38.32 22.44
CoBEVT(13, ) 27.09 48.47 66.82 55.20 40.93 25.33
Where2com(17, ) 28.53 46.57 66.17 55.42 41.92 23.35
SOAT methods. In this subsection, we used 9 SOAT cooperative perception algorithms (as shown in Subsection 5.1) to verify the performance of the 3D object tracking task in the U2UData dataset, including No Fusion, When2Com(16, ), DiscoNet(15, ), V2VNet(12, ), V2X-ViT(14, ), CoBEVT(13, ), and Where2com(17, ).
Evaluation. We utilize the same evaluation metrics as outlined in (tracking, ) for object tracking. These metrics include: AMOTA, average multiobject tracking accuracy; AMOTP, average multiobject tracking precision; sAMOTA, scaled average multiobject tracking accuracy, which ensures a more linear representation across the entire [0, 1] range of significantly challenging tracking tasks; MOTA, multi object tracking accuracy; MT, mostly tracked trajectories; ML, mostly lost trajectories.
Baselines tracker. We’ve chosen the AB3Dmot tracker from (tracking, ) as our baseline. This tracker initially retrieves 3D object detections from a LiDAR point cloud. It subsequently integrates the 3D Kalman filter with the birth and death memory technique to guarantee efficient and resilient tracking performance. It attains state-of-the-art performance while maintaining the fastest speed.

### v3

@@ 5.2. Cooperative 3D Object Tracking
When2Com(Liu et al., 2020) 24.30 43.68 60.69 52.62 36.75 26.21
DiscoNet(Li et al., 2021) 24.58 44.71 62.79 55.70 39.81 23.89
V2VNet(Wang et al., 2020) 26.19 47.43 65.81 56.58 41.63 23.83
V2X-ViT(Xu et al., 2022b) 26.85 46.50 64.50 55.92 38.32 22.44
CoBEVT(Xu et al., 2022a) 27.09 48.47 66.82 55.20 40.93 25.33
Where2com(Hu et al., 2022) 28.53 46.57 66.17 55.42 41.92 23.35
SOAT methods. In this subsection, we used 9 SOAT cooperative perception algorithms (as shown in Subsection 5.1) to verify the performance of the 3D object tracking task in the U2UData dataset, including No Fusion, When2Com(Liu et al., 2020), DiscoNet(Li et al., 2021), V2VNet(Wang et al., 2020), V2X-ViT(Xu et al., 2022b), CoBEVT(Xu et al., 2022a), and Where2com(Hu et al., 2022).
Evaluation. We utilize the same evaluation metrics as outlined in (Weng et al., 2020) for object tracking. These metrics include: AMOTA, average multiobject tracking accuracy; AMOTP, average multiobject tracking precision; sAMOTA, scaled average multiobject tracking accuracy, which ensures a more linear representation across the entire [0, 1] range of significantly challenging tracking tasks; MOTA, multi object tracking accuracy; MT, mostly tracked trajectories; ML, mostly lost trajectories.
Baselines tracker. We’ve chosen the AB3Dmot tracker from (Weng et al., 2020) as our baseline. This tracker initially retrieves 3D object detections from a LiDAR point cloud. It subsequently integrates the 3D Kalman filter with the birth and death memory technique to guarantee efficient and resilient tracking performance. It attains state-of-the-art performance while maintaining the fastest speed.